OpenStudy (anonymous):

how do i solve it? easy medal!!!!!!!!!!!!!!!!!!!!!! State the slope and the y-intercept for the graph of the equation. y=x+1

OpenStudy (anonymous):

Do they give you numbers for either y or x?

OpenStudy (anonymous):

y = mx + b is slope intercept form m is slope b is y-intercept. so 1 is the slope, and 1 is the y-intercept Make sense?

OpenStudy (anonymous):

nope XD @MelissaHolmes

OpenStudy (anonymous):

I graphed it here:

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

@Nathan28 ?

OpenStudy (anonymous):

still confused @MelissaHolmes could you walk me through it?

OpenStudy (anonymous):

sorry, I was out... had to go to school... :( if you still need the help... you can determine the slope and the y-intercept of an equation by using slope intercept form: y = mx + b. M is the slope and b is the y=intercept. so... y = mx + b y = 1x + 1 m = 1 b = 1 thus... slope = 1 y-intercept = 1 I hope I was still helpful... sorry so so so sorry for getting back on late.... :(

random231 (random231):

ya mellisa is correct @Nathan28

OpenStudy (anonymous):

actually XD i gave the wrong problem its y=x-1 no wonder why i was confused @random231

random231 (random231):

then u shud be able to do this one!!